OpenStudy (anonymous):

Find the fifth root of 32(cos(4.88692191)+isin(4.88692191)) or in other words 32(cos(280 degrees)+isin(280 degrees))

OpenStudy (anonymous):

the fifth root of 32 is 2 then \(\frac{280}{5}=56\) so one answer is \[2\left(\cos(56)+i\sin(56)\right)\]